OpenStudy (anonymous):

write the equation of the circle in general form.Endpoints of diameter:(1,1)(-1,-1)

OpenStudy (anonymous):

(x-1)(x+1) +(y-1)(y+1)=0 or x^2-1+y^2-1=0 x^2+y^2=2

OpenStudy (anonymous):

if (1,1)(-1,-1) are the endpoints of the diameter, it means that the midle point of this segment is the center. By cimetry, center will be at (0,0). To find the radius of this circle find length of this segment and devide it by 2. You find out that the radius is r=sqrt2 so equation of the circle will be:\[(x-0)^2+(y-0)^2=\sqrt2^2\]or \[x^2+y^2=2\]

OpenStudy (anonymous):

thanks guys both are great answers !